OpenStudy (k12awesomeness):

Medal and Fan!!! Solve for n. 3/4 x n = 1 A. B. C. D.

umerlodhi (umerlodhi):

k12 rocks

OpenStudy (k12awesomeness):

A. 3/4 B. 4/3 C. 1 3/4 D. 1 4/3

rebeccaxhawaii (rebeccaxhawaii):

multiply by the reciprocal

rebeccaxhawaii (rebeccaxhawaii):

example

OpenStudy (study312):

3 ÷ 4 = ?

OpenStudy (jdoe0001):

\(\bf \cfrac{3}{4}\times n=1\impliedby \textit{hint, multiply by 4 both sides}\)

rebeccaxhawaii (rebeccaxhawaii):



umerlodhi (umerlodhi):

Simplify 34xn34xn to 3xn43xn4 3xn4=1 3xn4=1 2 Multiply both sides by 44 3xn=1×4 3xn=1×4 3 Simplify 1×41×4 to 44 3xn=4 3xn=4 4 Divide both sides by 33 xn=43 xn=43 5 Divide both sides by nn x=43n

rebeccaxhawaii (rebeccaxhawaii):

so just flip the fraction

OpenStudy (study312):

site your sources

rebeccaxhawaii (rebeccaxhawaii):

so which one is that?

OpenStudy (k12awesomeness):

ohhh! So in MY case, it would be B ??

umerlodhi (umerlodhi):

medal i explained it bud

rebeccaxhawaii (rebeccaxhawaii):

yes

rebeccaxhawaii (rebeccaxhawaii):

@umerlodhi you were incorrect